1. A skater increases her velocity from 2.0 m/sec to 10 m/sec in 3.0 sec. what is her acceleration?
2. A cart rolling down an incline for 5.0 sec. If the cart has a beginning speed of 2.0 m/sec and an ending speed of 22 m/sec, what is its acceleration?

Answer by Alan3354(69443)

1. A skater increases her velocity from 2.0 m/sec to 10 m/sec in 3.0 sec. what is her acceleration?
(10 - 2)/3 = 8/3 m/sec/sec
-------------------
2. A cart rolling down an incline for 5.0 sec. If the cart has a beginning speed of 2.0 m/sec and an ending speed of 22 m/sec, what is its acceleration?
(22 - 2)/5 = 4 m/sec/sec